. An investigator wants to estimate the mean systolic blood pressure in children with congenital heart disease who are between the ages of 3 and 5. How many children should be enrolled in the study? The investigator plans on using a 95% confidence interval (so Z=1.96) and wants a margin of error of 5 units. The standard deviation of systolic blood pressure is unknown, but the investigators conduct a literature search and find that the standard deviation of systolic blood pressures in children with other cardiac defects is between 15 and 20. To estimate the sample size, we consider the larger standard deviation in order to obtain the most conservative (largest) sample size. 


1
Expert's answer
2021-05-07T09:21:27-0400

"n=(\frac{z_\frac{\alpha}{2 }s} {E}) ^{2 } =(\frac{1.96\times20} {5}) ^{2 } =61.47\approx 62"